# Langevin equation with potential of mean force: The case of anchored bath

### v1

@@ Abstract
The potential of mean force (PMF) is an effective average potential acting on an open system, renormalized due to the interaction with the surrounding thermal bath. The PMF is defined for an equilibrium ensemble, and generally it is not clear how to use it when the system is out of equilibrium and described by a (generalized) Langevin equation. We study a model where the system is a single particle (so there are no complications related to internal forces) and a non-trivial PMF is due to the presence of on-site (anchor) potentials applied to the bath particles. We found that the PMF does not merely replace the external potential, but also makes the dissipation kernel and statistical properties of noise dependent on the system’s position. That dependence is determined by the internal bath and system-bath interactions and is a priori unknown. Therefore, in the general case the Langevin equation with the PMF is not closed and thus inoperable. However, for systems with linear forces the aforementioned dependence on the system’s position may be canceled. As an example, we consider a model where the bath is formed by the Klein-Gordon chain, i. e. a harmonic chain with on-site harmonic potentials. In that case, the generalized Langevin equation has the standard form with an external potential replaced by a quadratic PMF.
@@ I Introduction
As discussed in the next section, the PMF is defined as a quantity that characterizes the system in thermal equilibrium. Clearly, the PMF must also affect the non-equilibrium dynamics described by a Langevin equation, but it is often not obvious in what way. The simplest assumption is that the Langevin equation and the corresponding fluctuation-dissipation relation retain their standard form and the only required modification is the replacement of the bare system potential by the PMF. That approach, broadly adopted in empirical studies, received theoretical endorsement in Refs. Lange ; Kinjo , where the standard generalized Langevin equation with the PMF was derived from the underlying dynamics using the projection operator technique. However, those derivations were later criticized in Refs. Schilling ; Schilling_review , with the conclusion that for systems with a PMF the standard Langevin equation can be derived neither exactly nor using any well-controlled approximation. On the other hand, in Refs. Vroy ; Vroy2 it was argued that a Langevin equation actually can be recovered from the first principles, but only at the expense of the noise and the dissipation kernel being position-dependent and with a modified fluctuation-dissipation relation. That conclusion seems to corroborate the earlier results for a system of several Brownian particles DO . Using a different approach, a generalized Langevin equation with an additional dissipation force, independent of the system’s velocity, was suggested in Netz .
@@ II Review of PMF
For future references, let us re-write $\langle F_{\alpha}\rangle$, using Eqs. (13) and (19), in terms of $Z_{0}(X)$:
$\langle F_{\alpha}\rangle=\frac{1}{\beta}\,\frac{\partial}{\partial X_{\alpha}}\,\ln Z_{0}(X).$ (20)
@@ III Anchored bath model
In the next two sections we shall use the Mazur-Oppenheim projection operator technique MO ; DO ; KO to derive a Langevin equation with the PMF for the model described in this section. The derivation is convenient to present in two steps. The first step is to re-write the equation of motion for the system separating the systematic and noisy components of the total force on the system. The result is what can be called the pre-Langeving equation, see Eq. (75) below. That equation is exact but impractical because it involves the system’s variables in an implicit form. The second step, discussed in Section 5, is to make that dependence explicit by applying an appropriate perturbation technique.
@@ IV Pre-Langevin equation
Here we take into account that the distribution $\rho_{0}=Z_{0}^{-1}\exp(-\beta H_{0})$ depends on $X$ through both $H_{0}$ and $Z_{0}$. Replacing $\partial H_{0}/\partial X=-F$ and using Eq. (20) this relation can be presented as
$\beta\left\langle F\,F^{\dagger}(\tau)\right\rangle-\beta\,\langle F^{\dagger}(\tau)\rangle\langle F\rangle+\left\langle\frac{\partial}{\partial X}F^{\dagger}(\tau)\right\rangle=0.$ (69)
From this we obtain
$\left\langle\frac{\partial}{\partial X}F^{\dagger}(\tau)\right\rangle=-\beta\left\langle E\,F^{\dagger}(\tau)\right\rangle,$ (70)
where $E$ is a zero-centered random variable corresponding to $F$,
$E=\mathcal{Q}F=F-\langle F\rangle,\quad\langle E\rangle=0.$ (71)
Note that $E$ is the initial value of the projected force $F^{\dagger}(t)=e^{\mathcal{Q}Lt}\,\mathcal{Q}F$,
$E=F^{\dagger}(0),$ (72)
and one can present $F^{\dagger}(t)$ as
$F^{\dagger}(t)=e^{\mathcal{Q}Lt}E.$ (73)
Then from Eqs. (67), (70), and (72) one obtains
$\mathcal{P}L_{s}F^{\dagger}(\tau)=\left(\frac{\partial}{\partial P}-\frac{\beta\,P}{M}\right)\langle E\,F^{\dagger}(\tau)\rangle.$ (74)
@@
$\dot{P}(t)=-\nabla V_{*}[X(t)]+F^{\dagger}(t)+\int_{0}^{t}d\tau\,e^{L(t-\tau)}\,\left(\frac{\partial}{\partial P}-\frac{\beta\,P}{M}\right)\langle E\,F^{\dagger}(\tau)\rangle.$ (75)
Compared to the case of a passive bath, this equation has two visible and intuitively expected differences. First, the external potential $V_{ex}$ is replaced by the PMF $V_{*}=V_{ex}+\Delta V$. Second, the integral term of the equation involves the correlation $\langle EF^{\dagger}(\tau)\rangle$ rather than $\langle FF^{\dagger}(\tau)\rangle$. The equation is exact, but not yet operational because $F^{\dagger}(t)$ depends on the system’s variables. That dependence can be made explicit by expanding the projected force in powers of a small mass ratio parameter $\lambda$, $F^{\dagger}(t)=F_{0}(t)+\lambda\,F_{1}(t)+\cdots$. For a passive bath, in the leading (zeroth) order in $\lambda$ the projected force $F^{\dagger}(t)\approx F_{0}(t)$ does not depend on the system’s variables, and Eq. (75) is readily reduced to the standard Langevin equation. However, as we shall see in the next section, for a non-passive bath $F^{\dagger}(t)$ depends on the system’s position even in the leading approximation. That dependence is a priori unknown, which may undermine the derivation of a practically useful Langevin equation with a PMF.
@@ V Langevin equation
$\dot{P}_{*}(t)=-\lambda\,\nabla V_{*}[X(t)]+\lambda\,F^{\dagger}(t)+\lambda^{2}\int_{0}^{t}d\tau\,e^{L(t-\tau)}\,\left(\frac{\partial}{\partial P_{*}}-\frac{\beta\,P_{*}}{m}\right)\langle E\,F^{\dagger}(\tau)\rangle.$ (77)
@@
$F^{\dagger}(t)=e^{\mathcal{Q}Lt}E=e^{(L_{0}+\lambda\mathcal{Q}L_{1})t}\,E,$ (80)
@@
$F^{\dagger}(t)=E_{0}(t)+\lambda E_{1}(t)+\mathcal{O}(\lambda^{2}).$ (82)
@@
$E_{0}(t)=e^{L_{0}t}E,\qquad E_{1}(t)=\frac{1}{m}\int_{0}^{t}d\tau\,P_{*}(t-\tau)\,e^{L_{0}(t-\tau)}\mathcal{Q}\frac{\partial}{\partial X}\,E_{0}(\tau).$ (83)
Both terms are zero-centered,
$\langle E_{0}(t)\rangle=\langle E_{1}(t)\rangle=0.$ (84)
Substituting $F^{\dagger}(t)\approx E_{0}(t)$ into Eq. (77), we find that in the leading order in $\lambda$ the equation has the form
$\dot{P}_{*}(t)=-\lambda\,\nabla V_{*}[X(t)]+\lambda\,E_{0}(t)-\frac{\lambda^{2}\beta}{m}\int_{0}^{t}d\tau\,P_{*}(t-\tau)\,\,e^{L(t-\tau)}\langle E\,E_{0}(\tau)\rangle.$ (85)
Returning from the scaled momentum $P_{*}=\lambda P$ to the true momentum $P$, we obtain
$\dot{P}(t)=-\nabla V_{*}[X(t)]+E_{0}(t)-\frac{\beta}{M}\int_{0}^{t}d\tau\,P(t-\tau)\,\,e^{L(t-\tau)}\langle E\,E_{0}(\tau)\rangle.$ (86)
The force $E_{0}(t)$ can be written as
@@
and interpreted as the deviation of the force, exerted by the bath on the system fixed at $X$, from the average value $\langle F(X)\rangle$. Clearly, $E_{0}(t)$ does not depend on $P$, but it does depend on $X$.
That dependence does not complicate the matter if the bath is a homogeneous non-anchored fluid. In that case, the dependence of forces on positions occurs only through the differences $X-x_{i}$ and $x_{i}-x_{j}$. When taking average over the bath coordinates $x_{i}$, one can change the integration variables $x_{i}^{\prime}=x_{i}-X$, so that the integrands depend on $x_{i}^{\prime}$ and $x_{i}^{\prime}-x_{j}^{\prime}$, but not on $X$. Therefore, although $F$ and $F_{0}(t)$ depend on $X$, their moments and correlations do not. As a result, $\langle F\rangle=0$ by symmetry, the PMF is reduced to the external potential, $\nabla V_{*}=\nabla V_{ex}$, and the correlation $\langle E\,E_{0}(t)\rangle=\langle F\,F_{0}(t)\rangle$ does not depend on $X$. Clearly, that correlation also does not depend on $P$ and bath variables. Therefore, the propagator $e^{Lt}$ in the integral term of Eq. (86) can be dropped, and the equation takes the form of the standard generalized Langevin equation
$\dot{P}(t)=-\nabla V_{ex}[X(t)]+F_{0}(t)-\frac{\beta}{M}\int_{0}^{t}d\tau\,P(t-\tau)\,\langle F\,F_{0}(\tau)\rangle$ (88)
with $F_{0}(t)$ as zero-centered noise, whose statistics do not depend on the system’s variables.
For an anchored bath, the situation is different because the statistical properties of noise $E_{0}(t)$ generally depend on $X$. While forces $F$ and $F_{0}(t)$ still depend on $X$ only through $X-x_{i}$, the internal bath forces depend not only on coordinate differences $x_{i}-x_{j}$, but also on the positions of individual bath units $x_{i}$. In that case, the variable change $x_{i}^{\prime}=x_{i}-X$ does not eliminate $X$ from the averages, so that $\langle F\rangle$ and $\langle FF_{0}(t)\rangle$ depend on $X$ and so does the correlation
$\langle EE_{0}(t)\rangle=\langle FF_{0}(t)\rangle-\langle F\rangle^{2}.$ (89)
@@
$\frac{\beta}{M}\langle EE_{0}(t)\rangle=\Gamma(t,X).$ (90)
@@
$\dot{P}(t)=-\nabla V_{*}[X(t)]-E_{0}(t)-\int_{0}^{t}d\tau\,P(t-\tau)\,\Gamma[\tau,X(t-\tau)],$ (91)
which formally is similar to the standard generalized Langevin equation (with the external potential replaced by the PMF), except that the dissipation term is not of the conventional convolution form since it depends on both $\tau$ and $t-\tau$. As it is, the equation is not operational because the dependence of the kernel on $X$ is a priory unknown. That is not necessarily a problem. After all, the exact form of the PMF $V_{*}(X)$ is usually not known either; it is often defined empirically to fit experimental data. One may define the function $\Gamma(X)$ empirically as well. However, the difficulty is that the two functions $V_{*}(X)$ and $\Gamma(X)$ are clearly not independent, yet there seems to be no general relation describing their connection.
The situation is simplified when the Hamiltonian $H_{0}$ has a special structure (34), $H_{0}=h_{1}(x,X)+h_{2}(X)$, and the force $F_{0}(t)$ on the fixed system has the form
$F_{0}(t)=f_{1}(X,x,t)+f_{2}(X).$ (92)
The component $f_{2}(X)$ depends only on the system’s position, while the component $f_{1}$ has the properties of the force exerted on a single-particle system by the homogeneous unanchored fluid, i. e. as in the standard theory of Brownian motion. Namely, the average of $f_{1}$ over bath variables (with distribution $\rho_{0}$) is zero for any $t$,
@@
and the time correlation
$\langle f_{1}(X,x,0)\,f_{1}(X,x,t)\rangle=c(t)$ (94)
does not depend on the system’s position $X$. In that case, the dependence of $\langle E\,E_{0}(t)\rangle$ on $X$ is canceled,
$\langle EE_{0}(t)\rangle=\langle F(0)\,F_{0}(t)\rangle-\langle F(0)\rangle^{2}=c(t)-c(0).$ (95)
Then in Eq. (86) the propagator $e^{L(t-\tau)}$ has no effect and can be dropped, and the equation turns into the generalized Langevin equation of the standard form
$\dot{P}(t)=-\nabla V_{ex}[X(t)]+F_{0}(t)-\frac{\beta}{M}\int_{0}^{t}d\tau\,P(t-\tau)\,\langle E\,E_{0}(\tau)\rangle$ (96)
with zero-centered noise $E_{0}(t)$ and the external potential $V_{ex}(X)$ replaced by the PMF $V_{*}(X)$. This is the equation advocated in Refs. Lange ; Kinjo . We observe that the equation is valid only under special assumptions (92)-(94). In the next two sections we consider a special (linear) version of the model with an anchored bath where assumptions (92) do hold and therefore the Langevin equation (96) is valid.

### v2

@@ Abstract
The potential of mean force (PMF) is an effective average potential acting on an open system, renormalized due to the interaction with the surrounding thermal bath. The PMF determines the correction to the equilibrium Gibbs distribution, but it is generally unclear how to implement the concept for the time-dependent phenomena described by a (generalized) Langevin equation. We study a model where the system is a single particle (so there are no complications related to internal forces) and a non-trivial PMF is due to the presence of on-site (anchor) potentials applied to the bath particles. We found that the PMF does not merely replace the external potential, but also makes the dissipation kernel and statistical properties of noise dependent on the system’s position. That dependence is determined by the internal bath and system-bath interactions and is a priori unknown. Therefore, in the general case the Langevin equation with the PMF is not closed and thus inoperable. However, for systems with linear forces the aforementioned dependence on the system’s position is canceled. As an example, we consider a model where the bath is formed by the Klein-Gordon chain, i. e. a harmonic chain with on-site harmonic potentials. In that case, the generalized Langevin equation has the standard form with an external potential replaced by a quadratic PMF.
@@ I Introduction
As discussed in the next section, the PMF is defined as a quantity that characterizes the equilibrium distribution and determines average equilibrium properties of the system. Clearly, the PMF must also affect time-dependent properties, such as time correlation and response functions, described by a Langevin equation - but how it affects them is not always obvious. The simplest assumption is that the Langevin equation and the corresponding fluctuation-dissipation relation retain their standard form and the only required modification is the replacement of the bare system potential by the PMF. That approach, broadly adopted in empirical studies, received theoretical endorsement in Refs. Lange ; Kinjo , where the standard generalized Langevin equation with the PMF was derived from the underlying dynamics using the projection operator technique. However, those derivations were later criticized in Refs. Schilling ; Schilling_review , with the conclusion that for systems with a PMF the standard Langevin equation can be derived neither exactly nor using any well-controlled approximation. On the other hand, in Refs. Vroy ; Vroy2 ; Lee it was argued that a Langevin equation actually can be recovered from the first principles, but only at the expense of the noise and the dissipation kernel being position-dependent and with a modified fluctuation-dissipation relation. That conclusion seems to corroborate the earlier results for a system of several Brownian particles DO . Using a different approach, a generalized Langevin equation with an additional dissipation force, independent of the system’s velocity, was suggested in Netz .
@@ III Anchored bath model
In the next two sections we shall use the Mazur-Oppenheim projection operator technique MO ; DO ; KO to derive a Langevin equation with the PMF for the model described in this section. The derivation is convenient to present in two steps. The first step is to re-write the equation of motion for the system separating the systematic and noisy components of the total force on the system. The result is what can be called the pre-Langevin equation, see Eq. (70) below. That equation is exact but impractical because it involves the system’s variables in an implicit form. The second step, discussed in Section 5, is to make that dependence explicit by applying an appropriate perturbation technique.
@@ IV Pre-Langevin equation
Here we take into account that for a non-passive bath the distribution $\rho_{0}=Z_{0}^{-1}\exp(-\beta H_{0})$ depends on $X$ through both $H_{0}$ and $Z_{0}$. However, the term with $\frac{\partial}{\partial X}\ln Z_{0}$ vanishes because of the factor $\langle F^{\dagger}(\tau)\rangle=0$. As a result, one obtains from Eq. (67)
$\left\langle\frac{\partial}{\partial X}F^{\dagger}(\tau)\right\rangle=-\beta\left\langle F\,F^{\dagger}(\tau)\right\rangle,$ (68)
and Eq. (66) takes the form
$\mathcal{P}L_{s}F^{\dagger}(\tau)=\left(\frac{\partial}{\partial P}-\frac{\beta\,P}{M}\right)\langle F\,F^{\dagger}(\tau)\rangle.$ (69)
@@
$\dot{P}(t)=-\nabla V_{*}[X(t)]+F^{\dagger}(t)+\int_{0}^{t}d\tau\,e^{L(t-\tau)}\,\left(\frac{\partial}{\partial P}-\frac{\beta\,P}{M}\right)\langle F\,F^{\dagger}(\tau)\rangle.$ (70)
This equation is similar to that for a passive bath MO ; KO , with the only visible difference that the external potential $V_{ex}$ is replaced by the PMF $V_{*}=V_{ex}+\Delta V$. A hidden difference is that the projected force $F^{\dagger}(t)=e^{\mathcal{Q}Lt}\mathcal{Q}F$ has the additional factor $\mathcal{Q}$, which ensures that $\langle F^{\dagger}(t)\rangle\sim\mathcal{PQ}=0$ for any time. The equation is exact, but not yet operational because $F^{\dagger}(t)$ depends on the system’s variables. That dependence can be made explicit by expanding the projected force in powers of a small mass ratio parameter $\lambda$, $F^{\dagger}(t)=F_{0}(t)+\lambda\,F_{1}(t)+\cdots$. For a passive bath, in the leading (zeroth) order in $\lambda$ the projected force $F^{\dagger}(t)\approx F_{0}(t)$ does not depend on the system’s variables, and Eq. (70) is readily reduced to the standard generalized Langevin equation. However, as we shall see in the next section, for a non-passive bath $F^{\dagger}(t)$ depends on the system’s position even in the leading approximation. That dependence is a priori unknown, which may undermine the derivation of a practically useful Langevin equation with a PMF.
@@ V Langevin equation
$\dot{P}_{*}(t)=-\lambda\,\nabla V_{*}[X(t)]+\lambda\,F^{\dagger}(t)+\lambda^{2}\int_{0}^{t}d\tau\,e^{L(t-\tau)}\,\left(\frac{\partial}{\partial P_{*}}-\frac{\beta\,P_{*}}{m}\right)\langle F\,F^{\dagger}(\tau)\rangle.$ (72)
@@
$F^{\dagger}(t)=e^{\mathcal{Q}Lt}\mathcal{Q}F=e^{(L_{0}+\lambda\mathcal{Q}L_{1})t}\,\mathcal{Q}F,$ (75)
@@
$F^{\dagger}(t)=E_{0}(t)+\lambda E_{1}(t)+\mathcal{O}(\lambda^{2}),$ (77)
@@
$E_{0}(t)=e^{L_{0}t}\mathcal{Q}F,\qquad E_{1}(t)=\frac{1}{m}\int_{0}^{t}d\tau\,P_{*}(t-\tau)\,e^{L_{0}(t-\tau)}\mathcal{Q}\frac{\partial}{\partial X}\,E_{0}(\tau),$ (78)
with both terms being zero-centered,
$\langle E_{0}(t)\rangle=\langle E_{1}(t)\rangle\sim\mathcal{PQ}=0.$ (79)
To order $\lambda^{2}$ the integral term in Eq. (72) is simplified and the equation takes the form
$\dot{P}_{*}(t)=-\lambda\,\nabla V_{*}[X(t)]+\lambda\,F^{\dagger}(t)-\frac{\lambda^{2}\beta}{m}\int_{0}^{t}d\tau\,P_{*}(t-\tau)\,\,e^{L(t-\tau)}\langle F\,E_{0}(\tau)\rangle.$ (80)
Since this equation is an approximation to order $\lambda^{2}$, it can only be used to evaluate properties which are either linear or quadratic in noise $\lambda\,F^{\dagger}(t)$. Upon averaging, the contributions linear in noise vanish, while contributions quadratic in noise are determined to order $\lambda^{2}$ by the leading term $E_{0}(t)$ in the expansion (77). Therefore, perturbation consistency requires to replace in Eq. (80) the projected force $F^{\dagger}(t)$ by the leading-order approximation $E_{0}(t)$. After that replacement, and after returning from the scaled to the true momentum $P=P_{*}/\lambda$, Eq. (81) takes the form
$\dot{P}(t)=-\nabla V_{*}[X(t)]+E_{0}(t)-\frac{\beta}{M}\int_{0}^{t}d\tau\,P(t-\tau)\,\,e^{L(t-\tau)}\langle F\,E_{0}(\tau)\rangle.$ (81)
The fluctuating force $E_{0}(t)$ can be written as
@@
and interpreted as the deviation of the force $F_{0}(t)$, exerted by the bath on the system fixed at $X$, from the average value $\langle F(X)\rangle$. Clearly, $E_{0}(t)$ does not depend on $P$, but it does depend on the system’s position $X$.
That dependence does not complicate the matter if the bath is a homogeneous non-anchored fluid. In that case, the dependence of forces on positions occurs only through the differences $X-x_{i}$ and $x_{i}-x_{j}$. When taking average over the bath coordinates $x_{i}$, one can change the integration variables $x_{i}^{\prime}=x_{i}-X$, so that the integrands depend on $x_{i}^{\prime}$ and $x_{i}^{\prime}-x_{j}^{\prime}$, but not on $X$. Therefore, although $F$ and $F_{0}(t)$ depend on $X$, their moments and correlations do not. Moreover, $\langle F\rangle=0$ by symmetry. As a result, the PMF is reduced to the external potential, $\nabla V_{*}=\nabla V_{ex}$, $E_{0}(t)$ coincides with $F_{0}(t)$, and the correlation $\langle F\,E_{0}(t)\rangle=\langle F\,F_{0}(t)\rangle$ does not depend on $X$. Clearly, that correlation also does not depend on $P$ and bath variables. Therefore, the propagator $e^{Lt}$ in the integral term of Eq. (81) can be dropped, and the equation takes the form of the standard generalized Langevin equation
$\dot{P}(t)=-\nabla V_{ex}[X(t)]+F_{0}(t)-\frac{\beta}{M}\int_{0}^{t}d\tau\,P(t-\tau)\,\langle F\,F_{0}(\tau)\rangle,$ (83)
supplemented with the ordinary fluctuation-dissipation relation.
For an anchored bath, the situation is more complicated because the statistical properties of noise $E_{0}(t)$ generally depend on $X$. While forces $F$ and $F_{0}(t)$ still depend on $X$ only through $X-x_{i}$, the internal bath forces depend not only on coordinate differences $x_{i}-x_{j}$, but also on the positions of individual bath units $x_{i}$. In that case, the variable change $x_{i}^{\prime}=x_{i}-X$ does not eliminate $X$ from the averages. As a result, $\langle F\rangle$ and $\langle FF_{0}(t)\rangle$ depend on $X$ and so does the correlation
$\langle FE_{0}(t)\rangle=\langle FF_{0}(t)\rangle-\langle F\rangle^{2}.$ (84)
@@
$\frac{\beta}{M}\langle FE_{0}(t)\rangle=\Gamma(t,X).$ (85)
@@
$\dot{P}(t)=-\nabla V_{*}[X(t)]-E_{0}(t)(t)-\int_{0}^{t}d\tau\,P(t-\tau)\,\Gamma[\tau,X(t-\tau)],$ (86)
which is formally similar to the standard generalized Langevin equation (with the external potential replaced by the PMF), except that the dissipation term is not of the conventional convolution form since it depends on both $\tau$ and $t-\tau$. As it is, the equation is not operational because the dependence of the kernel on $X$ is a priory unknown. That is not necessarily a problem. After all, the exact form of the PMF $V_{*}(X)$ is usually not known either; it is often defined empirically to fit experimental data. One may define the function $\Gamma(X)$ empirically as well. However, the difficulty is that the two functions $V_{*}(X)$ and $\Gamma(X)$ are clearly not independent, yet there seems to be no general relation describing their connection.
The situation is simplified when the Hamiltonian $H_{0}$ has a special structure (33), $H_{0}=h_{1}(x,X)+h_{2}(X)$, and therefore the force $F_{0}(t)$ on the fixed system has the form
$F_{0}(t)=e^{L_{0}t}\left(-\frac{\partial H_{0}}{\partial X}\right)=f_{1}(X,x,t)+f_{2}(X),$ (87)
where component $f_{2}(X)$ depends only on the system’s position. In addition, let us assume that component $f_{1}$ has the properties of noise in the standard theory of Brownian motion. Namely, the average of $f_{1}$ over bath variables (with distribution $\rho_{0}$) is zero for any $t$,
@@
and its auto-correlation correlation
$\langle f_{1}(X,x,0)\,f_{1}(X,x,t)\rangle\equiv\langle f_{1}f_{1}(t)\rangle=c(t)$ (89)
does not depend on the system’s position $X$. In that case,
$\langle F\rangle=f_{2}(X),\quad E_{0}(t)=F_{0}(t)-\langle F\rangle=f_{1}(X,x,t),$ (90)
and the correlation $\langle F\,E_{0}(t)\rangle$ coincides with the auto-correlation of $f_{1}$
$\langle FE_{0}(t)\rangle=\langle(f_{1}+f_{2})\,f_{1}(t)\rangle=\langle f_{1}f_{1}(t)\rangle+f_{2}\langle f_{1}(t)\rangle=\langle f_{1}f_{1}(t)\rangle,$ (91)
and therefore, because of assumption (89), does not depend on $X$. Then in Eq. (80) the propagator $e^{L(t-\tau)}$ has no effect and can be dropped. Also, since noise $E_{0}(t)$ is zero-centered, the correlation $\langle FE_{0}(t)\rangle$ coincides with the auto-correlation of $E_{0}(t)$,
$\langle E_{0}(0)\,E_{0}(t)\rangle=\langle F\,E_{0}(t)\rangle-\langle F\rangle\,\langle E_{0}(t)\rangle=\langle F\,E_{0}(t)\rangle.$ (92)
As a result, Eq. (80), written in terms of the true momentum $P=P_{*}/\lambda$, turns into the standard generalized Langevin equation
$\dot{P}(t)=-\nabla V_{ex}[X(t)]+E_{0}(t)-\frac{\beta}{M}\int_{0}^{t}d\tau\,P(t-\tau)\,\langle E_{0}(0)\,E_{0}(\tau)\rangle$ (93)
with the external potential $V_{ex}(X)$ replaced by the PMF $V_{*}(X)$. This is the equation advocated in Refs. Lange ; Kinjo . We observe that the equation is valid only under special assumptions. In the next two sections we consider a special (linear) version of the model with an anchored bath where assumptions (87)-(89) do hold and therefore the Langevin equation (93) is valid.